OpenStudy (anonymous):

A popular new spring hockey game that Jay got for his birthday uses springs to move a 0.0030-kg puck. It works like a regular table hockey game but instead of hitting the puck, the players use small springs. Each spring has a 120-N/m spring constant and can be compressed up to 0.020 m. How fast does the puck move when it departs a spring that was initially compressed this distance? If the spring is able to push the puck away in 0.2 seconds, what is the power generated by the spring?

OpenStudy (anonymous):

im not exactly sure which equations i should be using

OpenStudy (anonymous):

1/2kx^2=1/2mv^2 k=120N/m x=.02m m=.003Kg v=? 1/2*120*.02*.02=1/2*.003*v^2 v=4m/s p=work/t Work=1/2kx^2 =1/2*120*.02*.02=.024 J Power=.024/.2=.12w